Arturo's age is two years more than four times Allen's age. How old are Allen and Arturo respectively, if the sum of their ages is 32?

Answer by checkley79(3341)

X=(4Y+2)
X+Y=32
(4Y+2)+Y=32
5Y=32-2
5Y=30
Y=30/5
Y=6 YEARS OLD FOR ARTURO.
X=4*6+2=24+2=26 YEARS OLD FOR ALLEN.
PROOF:
26+6=32
32=32